OpenStudy (anonymous):

A daring 510-N swimmer dives off a cliff with a running horizontal leap,what must her minimum speed be just as she leaves the top of the cliff so that she will miss the ledge of the bottom, which is 1.75m wide and 9.00m below the top of the cliff.

OpenStudy (sidsiddhartha):

use simple kinematics

OpenStudy (sidsiddhartha):

ok are u familiar with this formula \[s=ut+\frac{ 1 }{ 2 }g t^2\]

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

whats the first step sir ??

OpenStudy (sidsiddhartha):

so here initial velocity that is u is equal to zero so u'll have \[s=\frac{ 1 }{ 2 }g t^2\] s=9.00 m that is given so \[9=\frac{ 1 }{ 2 }g t^2\rightarrow t=\sqrt{18/g}\] so \[t=1.35\] okay so far?

OpenStudy (anonymous):

yes i understand now and next ?

OpenStudy (abhisar):



OpenStudy (anonymous):

by step sir please so i can practice

OpenStudy (sidsiddhartha):

now it is given that she leaves the top of the cliff so that she will miss the ledge of the bottom, which is 1.75m wide so she must travel 1.75 meter now simply use \[d=(vecocity)*(time)=v*t\] so \[v=\frac{ d }{ t }=\frac{ 1.75 }{ 1.35 }=1.29 m/s\]

OpenStudy (sidsiddhartha):

got this?

OpenStudy (abhisar):

@sidsiddhartha xplained it very well but just to conclude

OpenStudy (sidsiddhartha):

yeah a diagram is always better :)

OpenStudy (anonymous):

yup @sidsiddhartha

OpenStudy (abhisar):

Step1 Find the time taken in travelling 9m from the top of the cliff. Suppose u got t seconds Step2 Find the velocity required to travel 1.75 m in t seconds

OpenStudy (anonymous):

@Abhisar i know that sir but i didnt know the formula already

OpenStudy (abhisar):

@sidsiddhartha explained everything very nicely....He even did the calculations for you. I hope you know the formulas now !

OpenStudy (abhisar):

do u ?

OpenStudy (anonymous):

only that is the formula ???

OpenStudy (abhisar):

There is no direct formula for this question (if u r expecting that). Two formulas are used here, \(\large\sf S=ut+\frac{1}{2}gt^2\) and Speed = Distance/time

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

sir can you teel the unknown

OpenStudy (abhisar):

s=distance or height [9 m in this case] u=initial speed [0 in this case] t=time taken [1.35 seconds] g=acceleration due to gravity [9.8 m/s]

OpenStudy (anonymous):

so what 1.29 m/s ?? is that the initial velocity ??

OpenStudy (abhisar):

It is horizontal speed by which the diver should run in order to just miss the ledge.

OpenStudy (abhisar):



OpenStudy (anonymous):

so thats the ans. ???????????

OpenStudy (abhisar):

0 is the vertical component of initial velocity and 1.29 is the horizontal component of initial velocity. Yes the answer is 1.29 m/s

OpenStudy (abhisar):

Have you learned projectile motion ?

OpenStudy (anonymous):

not yet sir but this my ps in projectile motion

OpenStudy (abhisar):

\(\underline{\href{//www.physicsclassroom.com/class/vectors/u3l2a.cfm}{\sf \huge This}}\) can be helpful in your learning

OpenStudy (anonymous):

thank you sir ,, sir explain everything plsssss

OpenStudy (abhisar):

i would hav explained you but it is very long topic and i doubt i have that much time ryt now. I'll urge you to go through it by urself. Mark the doubts and ask me next time. Can you do this ?

OpenStudy (anonymous):

mm can you solve it to me sir

OpenStudy (anonymous):

sir

OpenStudy (abhisar):

Solve what ?

OpenStudy (anonymous):

this i post si

OpenStudy (abhisar):

but it's already solved !

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @sidsiddhartha so here initial velocity that is u is equal to zero so u'll have \[s=\frac{ 1 }{ 2 }g t^2\] s=9.00 m that is given so \[9=\frac{ 1 }{ 2 }g t^2\rightarrow t=\sqrt{18/g}\] so \[t=1.35\] okay so far? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @sidsiddhartha now it is given that she leaves the top of the cliff so that she will miss the ledge of the bottom, which is 1.75m wide so she must travel 1.75 meter now simply use \[d=(vecocity)*(time)=v*t\] so \[v=\frac{ d }{ t }=\frac{ 1.75 }{ 1.35 }=1.29 m/s\] \(\color{blue}{\text{End of Quote}}\)